Calculate the amount of heat required to heat a 3.1 kg gold bar from 25 ∘C to 69 ∘C. Specific heat capacity of gold is 0.128 J/g∘C
1
Expert's answer
2017-01-30T14:45:09-0500
Q = c*m*ΔT = 0.128*3100*(69-25) = 17459 (J) = 17.5 (kJ)